this is the qustion solve for x in terms of A and P heres the equation A=4P(1+x)^2 i have no idea where to start the square throws me off.
Answer by Alan3354(69443)

solve for x in terms of A and P heres the equation A=4P(1+x)^2
----------------
A=4P(1+x)^2
(1+x)^2 = A/4P
1+x = sqrt(A/4P) = sqrt(AP)/2P
x = -1 +/- (sqrt(AP)/2P)